OpenStudy (anonymous):

If f(x)=(x+1)^−1 and g(x)=x−2, what is the domain of f(x)÷g(x)? A) all values of x B) (−∞,−1),(−1,2), and (2,∞) C) (−∞,2) and (2,∞) D) (−∞,−1] and [2,∞) i don't quite understand how to get the answer. help plz.

OpenStudy (anonymous):

first set up f(x)/g(x)

OpenStudy (anonymous):

f(x)=1/(1+x) /(g(x) dividing by g(x) is the same as times by reciprocal .

OpenStudy (anonymous):

so instead of (x+1)^-1/ x-2, it would be x-2*(x+1)^-1

OpenStudy (anonymous):

Yes, 1/ploynomial means there will be an asymtote

OpenStudy (solomonzelman):

well, \(\large\color{black}{ \displaystyle f(x)=(x+1)^{-1} }\) \(\large\color{black}{ \displaystyle g(x)=x-2 }\) then, \(\large\color{black}{ \displaystyle \frac{f(x)}{g(x)}~~~~\Rightarrow~~~~\frac{(x+1)^{-1}}{x-2} }\)

OpenStudy (solomonzelman):

but that is then same as \(\large\color{black}{ \displaystyle \frac{f}{g}(x)~~=~~\frac{1}{(x-1)(x-2)} }\)

OpenStudy (solomonzelman):

so there are 2 values at which the function is undefined ... what are they?

OpenStudy (anonymous):

in a factored equation like (x+1)*(x-2) the places where it would cross the x axis would be -1 and +2. when you do 1/(poly) , those points become the asymtotes.